OpenStudy (anonymous):

A total of 801 tickets were sold for the school play. They were either adult tickets or student tickets. There were 51 more student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (anonymous):

a + s = 801 s=a+51 a + a + 51 = 801 2a + 51=801 2a = 750 a=375